OpenStudy (anonymous):

Medal and fan!!!!!!! Evaluate each expression for the given value of the variable. -2a^-1 b^-3 for a=3 and b=3.

OpenStudy (mathmale):

Any chance you could try this problem yourself first, and then share the results, before offering medals and fan status? Given the expression 2a, substitute a=3. What do you get? Hope that puts you onto the right track.

OpenStudy (anonymous):

sure mathmale i already have my answer just checking if its right. i got -2/54

OpenStudy (mathmale):

Given -2a^(-1), substitute a=3. What do you get? Given b^(-3), substitute b=3. What do you get?

OpenStudy (anonymous):

since the exponents of the variables are negative can i do this -2/a*b^3

OpenStudy (mathmale):

Not if I read your math problem statement correctly. As I read it, you are given two separate math expressions involving exponents. You are to evaluate them separately: Given -2a^(-1), substitute a=3. What do you get? Given b^(-3), substitute b=3. What do you get?

OpenStudy (anonymous):

no its actually (-2a^-1)*(b^(-3))

OpenStudy (anonymous):

so just one expression

OpenStudy (mathmale):

OK. Thanks. Then the first factor is -2*3^(-1) = -2/3 and the second factor is 3^(-3) = 1/27 and you need to multiply those two fractions together to obtain the final desired result.